# ContrastScore: Towards Higher Quality, Less Biased, More Efficient Evaluation Metrics with Contrastive Evaluation

### v1

@@ 4 Experimental Setup
We evaluate the performance and robustness of the proposed ContrastScore on two tasks: machine translation and summarization, using models from the Llama and Qwen families to ensure generalizability. To benchmark ContrastScore, we compare it against single-model evaluators as well as an ensemble setting, where the probabilities of the expert and amateur models are averaged per token. The ensemble probability at token $t$, denoted as $p_{\text{Ens}}^{t}$, is computed as: $p_{\text{Ens}}^{t}=\frac{1}{2}\left(p_{\text{EXP}}^{t}+p_{\text{AMA}}^{t}\right)$.
@@ 4.1 Datasets and Tasks
For machine translation experiments, we utilize the WMT23 Metrics Shared Task dataset Freitag et al. (2023), focusing on three language pairs: English-German (EN-DE), Chinese-English (ZH-EN), and Hebrew-English (HE-EN). These pairs were selected to cover a diverse range of translation scenarios, including high-resource (EN-DE), typologically distant (ZH-EN), and low-resource (HE-EN) settings, ensuring a comprehensive evaluation of our approach. For summarization, we use the SummEval dataset Fabbri et al. (2021), which consists of summaries generated from CNN/DailyMail articles.
@@ 4.2 Meta-Evaluation
Correlation with Human Scores (Quality).  The effectiveness of an evaluator is measured using the Pearson correlation between its scores and human scores, where a higher correlation indicates stronger alignment with human judgments. For translation, Pearson correlation assesses the evaluator’s ability to capture differences between candidate translations based on the source texts, comparing them to human judgment scores derived from fine-grained MQM annotations (Freitag et al., 2021). For summarization, we compute Pearson correlation with human judgments across four quality aspects: coherence, fluency, consistency, and relevance.
Bias Evaluation.  We measure biases in two dimensions: likelihood bias and length bias.
Likelihood Bias: Recent studies have shown that LLMs tend to favor sentences they deem more likely, often assigning higher evaluation scores and performing better on such sentences, regardless of the specific task or evaluation criteria Ohi et al. (2024); McCoy et al. (2024a, b). Ohi et al. (2024) find that LLM-based evaluators systematically overrate high-likelihood sentences and underrate low-likelihood ones compared to human scores. To quantify this likelihood bias, they propose BiasScore, defined as:
@@
Length Bias: It is observed that summarization evaluation metrics tend to favor longer summaries, particularly in source-based settings Sun et al. (2019); Liu et al. (2024). Liu et al. (2024) find that LLM-based evaluators exhibit this length bias, often assigning higher scores to longer texts regardless of quality. Similarly, we quantify this tendency using the Spearman correlation between evaluator scores and the length of the generated text, where a lower correlation magnitude indicates a weaker length bias, suggesting a more content-focused evaluation.
Efficiency.  An evaluator’s efficiency is measured by the number of samples processed per second, with a higher processing speed indicating greater efficiency.
Source-based Reference-based
Llama EN-DE ZH-EN HE-EN Avg EN-DE ZH-EN HE-EN Avg
1B 0.530 0.518 0.320 0.456 0.608 0.499 0.456 0.521
Single 3B 0.578 0.526 0.293 0.466 0.628 0.498 0.466 0.531
8B 0.632 0.574 0.491 0.566 0.636 0.498 0.474 0.536
(3B,1B) 0.568 0.538 0.306 0.471 0.629 0.499 0.454 0.528
Ensemble (8B,1B) 0.605 0.575 0.462 0.547 0.639 0.505 0.461 0.535
(8B,3B) 0.619 0.574 0.453 0.549 0.645 0.505 0.466 0.539
(3B,1B) 0.599 0.562 0.284 0.482 0.643 0.532 0.463 0.546
ContrastScore (8B,1B) 0.641 0.590 0.482 0.571 0.653 0.541 0.484 0.559
(8B,3B) 0.639 0.595 0.482 0.572 0.652 0.542 0.486 0.560
(a) Correlation of Llama-based evaluators with human scores.
Source-based Reference-based
Qwen EN-DE ZH-EN HE-EN Avg EN-DE ZH-EN HE-EN Avg
0.5B 0.487 0.557 0.347 0.463 0.577 0.509 0.469 0.518
Single 3B 0.573 0.594 0.415 0.527 0.599 0.475 0.477 0.517
7B 0.600 0.574 0.445 0.540 0.608 0.480 0.478 0.522
(3B,0.5B) 0.548 0.599 0.421 0.523 0.610 0.503 0.473 0.528
Ensemble (7B,0.5B) 0.567 0.598 0.446 0.537 0.616 0.505 0.476 0.532
(7B,3B) 0.599 0.608 0.456 0.554 0.624 0.502 0.482 0.536
(3B,0.5B) 0.590 0.628 0.431 0.550 0.627 0.531 0.480 0.546
ContrastScore (7B,0.5B) 0.611 0.628 0.465 0.568 0.632 0.535 0.489 0.552
(7B,3B) 0.605 0.629 0.466 0.567 0.629 0.535 0.490 0.551
(b) Correlation of Qwen-based evaluators with human scores .
Table 1: Correlation of Llama and Qwen-based evaluators with human scores on the WMT23 dataset. Correlation values that are highest within each group (i.e. Single, Ensemble and ContrastScore) are underlined. Additionally, the highest correlation for each language pair is boldfaced.
@@ 4.3 Hyperparameters and Environment
To ensure robust generalization, we use models from the same family but of different sizes, where a larger model serves as the expert and a smaller model as the amateur. Specifically, we use LLama3.2-Instruct (1B, 3B) and LLama3.1-Instruct (8B), as the LLama 3.2 version does not include an 8B model. Additionally, we evaluate Qwen2.5-Instruct (0.5B, 3B, and 7B). We also attempted to explore ContrastScore using models from different families; however, this proved problematic due to differences in model tokenization. In all our experiments [^2] We set the decoding temperatures for both the expert and amateur models, denoted as $\tau_{1}$ and $\tau_{2}$, respectively. Following Li et al. (2023), we set $\tau_{2}=0.5$, while $\tau_{1}$ is set to 1.5., the scaling factor for the amateur model, $\gamma$, is set to 0.1. This setting is based on our pilot investigation, which found that $\gamma=0.1$ gives the best overall performance empirically for the ZH-EN language pair of the machine translation task, see Appendix B, Figure 2. All experiments were run on a GPU cluster under same conditions and node configuration to ensure fair comparisons. Each node of the cluster has 4$\times$×H100 GPUs.
@@ 5 Results
We assess the effectiveness of ContrastScore across multiple aspects, including its correlation with human scores, bias mitigation, and computational efficiency, in comparison to single-model and ensemble methods.
Llama Coherence Consistency Fluency Relevance Avg.
1B 0.417 0.549 0.542 0.344 0.463
Single 3B 0.388 0.575 0.530 0.338 0.458
8B 0.368 0.588 0.539 0.332 0.457
(3B,1B) 0.427 0.564 0.542 0.344 0.469
Ensemble (8B,1B) 0.429 0.581 0.556 0.357 0.481
(8B,3B) 0.414 0.587 0.544 0.351 0.474
(3B,1B) 0.456 0.554 0.551 0.334 0.474
ContrastScore (8B,1B) 0.461 0.584 0.575 0.368 0.497
(8B,3B) 0.462 0.587 0.575 0.369 0.498
(a) Correlation of Llama-based evaluators with human scores
Qwen Coherence Consistency Fluency Relevance Avg.
0.5B 0.386 0.536 0.520 0.315 0.439
Single 3B 0.370 0.544 0.492 0.319 0.431
7B 0.371 0.560 0.510 0.327 0.442
(3B,0.5B) 0.394 0.568 0.536 0.331 0.457
Ensemble (7B,0.5B) 0.386 0.577 0.538 0.333 0.459
(7B,3B) 0.378 0.584 0.521 0.330 0.453
(3B,0.5B) 0.397 0.594 0.546 0.323 0.465
ContrastScore (7B,0.5B) 0.376 0.608 0.554 0.322 0.465
(7B,3B) 0.384 0.606 0.557 0.334 0.470
(b) Correlation of Qwen2.5-based evaluators with human scores
Table 2: Correlation of Llama and Qwen-based evaluators with human scores on the SummEval dataset. Correlation values that are relatively highest within each group (i.e. Single, Ensemble and ContrastScore) are underlined. Additionally, the highest correlation for each evaluation dimension is boldfaced.
@@ 5.1 Correlation with Human Scores
Machine Translation.  Table 1 presents the Pearson correlation between various evaluators and human scores for both source-based and reference-based settings, following Qin et al. (2023). We additionally present a visual comparison in Fig. 4 in Appendix C. The results show that ContrastScore consistently outperforms both single-model and ensemble approaches for both settings. For example, ContrastScore based on Qwen(7B, 0.5B) improves correlation by 5.2% over the single Qwen 7B model and by 5.8% over the ensemble of Qwen(7B, 0.5B) in the source-based setting. It reveals that ensembling does not always improve performance. While ensemble methods are generally expected to refine predictions, they can sometimes reduce correlation with human scores. For instance, ensembling Llama (8B, 3B) in the source-based setting leads to a 3.0% decrease in correlation compared to the single Llama 8B model. Similarly, the ensemble of Llama (3B, 1B) in reference-based setting results in a marginal 0.6% reduction compared to the single Llama 3B model. Additionally, we observe that ContrastScore’s effectiveness depends on the reliability of the amateur model. When the amateur model fails to provide a meaningful correction signal, ContrastScore may not yield improvements. This is evident in the HE-EN pair for Llama models in the source-based setting, where the correlation scores for both the 1B and 3B models (0.32 and 0.29, respectively) are too low to serve as reliable correction signals. Consequently, ContrastScore does not surpass the single 8B model’s correlation in this case.
Beyond these cases where the expert model underperforms relative to the amateur model or when the amateur model is too weak to provide a meaningful correction signal, our results demonstrate that ContrastScore can achieve correlation scores comparable to, or even exceeding, those of the largest model—using only two smaller models. This has significant efficiency implications, as evaluation metrics based on LLMs are used frequently throughout the model development lifecycle, and inference costs scale with model size. Since larger models are computationally expensive, being able to obtain reliable evaluation with smaller models presents a practical advantage. This efficiency gain is particularly valuable for scalable and cost-effective evaluation pipelines. We further explore this trade-off in §5.3.
Summarization.   Table 2 presents the Pearson correlation between different evaluators and human scores for summarization. We additionally present a visual comparison in Fig. 5 in Appendix C. Similar to prior work Liu et al. (2023); Fu et al. (2024), we do not include reference-based evaluation for summarization, as source-based evaluation has been shown to be more robust. Reference-based metrics often struggle with the high variability of reference summaries, leading to weaker correlations with human judgments. In contrast, source-based approaches better assess factual consistency and relevance based on more informative sources, making them more reliable for evaluating summarization quality.
Our results indicate that ContrastScore provides best results overall. For instance, ContrastScore with Llama(8B, 3B) improves correlation by 9.0% over the single Llama 8B, whereas ensembling Llama(8B, 3B) only provides a 3.7% gain. This further supports the effectiveness of ContrastScore as a superior alternative to ensembling. Beyond overall performance, ContrastScore consistently surpasses single models while utilizing significantly fewer parameters. Specifically, ContrastScore with Qwen(3B, 0.5B) outperforms the single Qwen 7B by 5.2% , and ContrastScore with Llama(3B, 1B) exceeds the single Llama 8B by 3.7%. These results reinforce the observation that ContrastScore can efficiently enhance model evaluation by leveraging smaller models.
@@
Besides improving overall correlation with human scores, ContrastScore demonstrates notable gains across key summarization dimensions, particularly coherence, fluency, and relevance. When applied to Llama(8B, 3B), it enhances coherence by 25.5%, fluency by 6.7%, and relevance by 11.1% compared to the single Llama 8B. However, its impact on consistency differs across model families. While it consistently improves consistency scores in Qwen-based models, its effectiveness in Llama-based models is less predictable. This suggests that the way ContrastScore interacts with different evaluation dimensions may depend on model-specific characteristics.
Discussion.  Overall, ContrastScore consistently outperforms both single-model and ensemble approaches across diverse settings, demonstrating its robustness in aligning with human judgments. While it enhances evaluation performance across tasks, it is particularly effective in summarization, substantially improving coherence, fluency, and relevance. In machine translation, it also provides notable gains, especially for high-resource language pairs. Moreover, ContrastScore maintains strong evaluation performance while leveraging smaller models, making it a computationally efficient alternative to single-model methods.
@@ 5.2 Bias Analysis
Likelihood Bias. Table 3 presents the likelihood bias scores across different evaluators. The results clearly demonstrate that ContrastScore substantially mitigates likelihood bias compared to both single-model and ensemble approaches. Specifically, ContrastScore with Llama (8B, 3B) reduces the likelihood bias by 37.0% and 61.1% compared to single Llama 8B in summarization and machine translation task, respectively. Similarly, ContrastScore with Qwen (7B, 0.5B) results in a reduction of 25.6% and 34.8% compared to single Qwen 7B model for the same tasks.
Llama Summ WMT Qwen Summ WMT
Single 1B 0.348 0.212 0.5B 0.376 0.252
3B 0.385 0.245 3B 0.392 0.451
8B 0.381 0.352 7B 0.398 0.463
Ensemble (3B,1B) 0.308 0.123 (3B,0.5B) 0.379 0.327
(8B,1B) 0.326 0.152 (7B,0.5B) 0.393 0.294
(8B,3B) 0.359 0.229 (7B,3B) 0.369 0.381
Contrast (3B,1B) 0.183 0.026 (3B,0.5B) 0.236 0.314
(8B,1B) 0.262 0.134 (7B,0.5B) 0.296 0.302
(8B,3B) 0.240 0.137 (7B,3B) 0.329 0.287
Table 3: Likelihood bias scores for machine translation and summarization tasks across Qwen and Llama model families. The relatively highest bias value within each group is underlined. The lowest overall bias score is boldfaced.
Llama Summ WMT Qwen Summ WMT
Single 1B 0.218 -0.692 0.5B 0.289 -0.686
3B 0.289 -0.529 3B 0.391 -0.525
8B 0.344 -0.500 7B 0.408 -0.460
Ensemble (3B,1B) 0.212 -0.633 (3B,0.5B) 0.349 -0.590
(8B,1B) 0.274 -0.628 (7B,0.5B) 0.373 -0.562
(8B,3B) 0.301 -0.564 (7B,3B) 0.403 -0.505
Contrast (3B,1B) 0.086 -0.677 (3B,0.5B) 0.253 -0.599
(8B,1B) 0.206 -0.632 (7B,0.5B) 0.323 -0.565
(8B,3B) 0.201 -0.632 (7B,3B) 0.329 -0.579
Table 4: Length bias scores for machine translation and summarization tasks. The relatively highest bias value within each group is underlined. The lowest overall bias score is boldfaced.
Length Bias.  Table 4 shows the length bias scores for different models across summarization and machine translation tasks. We observe that ContrastScore considerably reduces length bias in summarization. Notably, ContrastScore with Llama (3B,1B) achieves a substantial reduction of 70.2%, lowering the bias score from 0.289 to 0.086. Similarly, for Qwen (3B,0.5B), ContrastScore reduces the bias score from 0.349 to 0.253, demonstrating its effectiveness in mitigating length bias in summarization. While the impact of length bias in machine translation is less well understood and has received comparatively limited attention, we include length bias scores in machine translation task for completeness. As expected, we do not observe a clear or consistent trend in the machine translation setting.

### v2

@@ 4.1 Datasets and Tasks
For summarization, we use the SummEval dataset Fabbri et al. (2021), which contains model-generated summaries of CNN/DailyMail articles, as well as QAGS-XSUM Wang et al. (2020), which includes 239 system outputs on the XSUM dataset. For machine translation, we use the MQM22 and MQM23 datasets from the WMT22 Freitag et al. (2022) and WMT23 Freitag et al. (2023) Metrics Shared Tasks. The language pairs involved represent a diverse range of translation scenarios, including high-resource (EN-DE), typologically distant (ZH-EN and EN-RU), and low-resource (HE-EN) settings, thereby enabling a comprehensive evaluation of our approach. See AppendixA.1 for more details.
4.2 Baseline Metrics
We consider the following baseline metrics for comparison: BLEU Papineni et al. (2002), CHRF Papineni et al. (2002), ROUGE 1, ROUGE 2, ROUGE L Lin (2004), COMET Rei et al. (2020)COMET-KIWI Rei et al. (2022), BERTScore Zhang et al. (2020), MoverScore Zhao et al. (2019), and BARTScore Yuan et al. (2021). The details of these metrics can be found in Appendix A.2. In addition, we compare it against Single-model evaluators as well as an Ensemble setting, where the probabilities of the expert and amateur models are averaged per token, using models from the LLaMA and Qwen families to ensure generalizability. The ensemble probability at token $t$, denoted as $p_{\text{Ens}}^{t}$, is computed as:
$p_{\text{Ens}}^{t}=\frac{1}{2}\left(p_{\text{EXP}}^{t}+p_{\text{AMA}}^{t}\right)$ (6)
@@ 4.3 Meta-Evaluation
Correlation with Human Scores (Quality).  The effectiveness of an evaluator is measured using the Pearson correlation between its scores and human scores, where a higher correlation indicates stronger alignment with human judgments. For translation, Pearson correlation assesses the evaluator’s ability to capture differences between candidate translations based on the source texts, comparing them to human judgment scores derived from fine-grained MQM annotations (Freitag et al., 2021). For summarization, we compute Pearson correlation with human judgments across different quality aspects: coherence, fluency, consistency, relevance, and factuality.
Bias Evaluation.  We measure biases in likelihood bias. Recent studies have shown that LLMs tend to favor sentences they deem more likely, often assigning higher evaluation scores and performing better on such sentences, regardless of the specific task or evaluation criteria Ohi et al. (2024); McCoy et al. (2024a, b). Ohi et al. (2024) find that LLM-based evaluators systematically overrate high-likelihood sentences and underrate low-likelihood ones compared to human scores. To quantify this likelihood bias, they propose BiasScore defined as:
@@
Efficiency.  It is measured by the number of samples processed per second, with a higher processing speed indicating greater efficiency.
@@ 4.4 Hyperparameters and Environment
[…]
@@ 5 Results
We evaluate ContrastScore in terms of its correlation with human scores, bias, and efficiency, compared to single-model and ensemble methods.
@@ 5.1 Correlation with Human Scores
Evaluators QAGS-XSUM SummEval AVG
Factuality Coherence Consistency Fluency Relevance
Baseline ROUGE-1 0.055 0.238 0.079 0.071 0.330 0.154
ROUGE-2 0.121 0.164 0.073 0.066 0.229 0.130
ROUGE-L 0.075 0.207 0.071 0.078 0.303 0.147
MoverScore 0.028 0.129 0.141 0.133 0.280 0.142
BERTScore 0.024 0.342 0.115 0.156 0.372 0.202
BARTScore 0.099 0.450 0.339 0.343 0.428 0.332
LLaMA Single 1B 0.064 0.417 0.549 0.542 0.344 0.383
3B 0.157 0.388 0.575 0.530 0.338 0.398
8B 0.205 0.368 0.588 0.539 0.332 0.406
Ensemble (3B,1B) 0.170 0.427 0.564 0.542 0.344 0.409
(8B,1B) 0.212 0.429 0.581 0.556 0.357 0.427
(8B,3B) 0.241 0.414 0.587 0.544 0.351 0.428
Contrast (3B,1B) 0.262 0.456 0.554 0.551 0.334 0.431
(8B,1B) 0.324 0.461 0.584 0.575 0.368 0.462
(8B,3B) 0.323 0.462 0.587 0.575 0.369 0.463
Qwen Single 0.5B 0.107 0.386 0.536 0.520 0.315 0.373
3B 0.112 0.370 0.544 0.492 0.319 0.367
7B 0.091 0.371 0.560 0.510 0.327 0.372
Ensemble (3B,0.5B) 0.081 0.394 0.568 0.536 0.331 0.382
(7B,0.5B) 0.095 0.386 0.577 0.538 0.333 0.386
(7B,3B) 0.103 0.378 0.584 0.521 0.330 0.383
Contrast (3B,0.5B) 0.128 0.397 0.594 0.546 0.323 0.398
(7B,0.5B) 0.122 0.376 0.608 0.554 0.322 0.396
(7B,3B) 0.134 0.384 0.606 0.557 0.334 0.403
Table 1: Pearson correlation of evaluators with human scores in summarization. boldface represents best overall scores, while underline represents best scores within each model group (Baseline, LLaMA, Qwen). Overall, ContratScore outperforms single and ensemble methods as well as baseline metrics for both LLaMA and Qwen families.
@@
Summarization.  Table 1 presents the Pearson correlation between different evaluators and human scores for summarization. Our results indicate that ContrastScore provides best results overall. Furthermore, ContrastScore consistently surpasses single models while utilizing much fewer parameters. Specifically, ContrastScore with Qwen(3B, 0.5B) outperforms the single Qwen 7B by 7.0%, and ContrastScore with LLaMA(3B, 1B) exceeds single LLaMA 8B by 6.2%. These results reinforce the observation that ContrastScore can efficiently enhance model evaluation by leveraging smaller models. Besides improving overall correlation with human scores, ContrastScore demonstrates notable gains across key summarization dimensions, particularly coherence, fluency, relevance and factuality. When applied to LLaMA(8B, 3B), it enhances coherence by 25.5%, fluency by 6.7%, relevance by 11.1%, and factuality by 57.6%, compared to the single LLaMA 8B. However, its impact on consistency differs across model families.
Machine Translation.  Table 2 presents the Pearson correlation between various evaluators and human scores. The results show that, overall, ContrastScore outperforms both single-model and ensemble approaches. For example, ContrastScore based on Qwen(7B, 3B) improves correlation by 7.1% over the single Qwen 7B model and by 4.6% over the ensemble of Qwen(7B, 0.5B). In contrast, ensembling does not always improve performance: While ensemble methods are generally expected to refine predictions, they can sometimes reduce correlation with human scores. For instance, ensembling LLaMA (8B, 1B) leads to a 3.3% decrease in correlation compared to the LLaMA 8B model. Similar trends can be observed in pairwise accuracy Deutsch et al. (2023), as shown in Table 10 of the Appendix.
Beyond the cases when the amateur model is too weak to provide a meaningful correction signal—such as the low correlation scores observed in the HE-EN pair for both LLaMA 1B and 3B (0.32 and 0.29, respectively), which fail to offer reliable contrastive feedback against LLaMA 8B—our results demonstrate that ContrastScore can achieve correlation scores comparable to, or even exceeding, those of the largest model using only two smaller models. Specifically, ContratScore of Qwen(3B, 0.5B) surpasses single Qwen 7B by 3.1%, improving from 0.449 to 0.463 with only half the parameters.
Evaluators MQM22 MQM23 AVG
EN-DE ZH-EN EN-RU EN-DE ZH-EN HE-EN
Baseline BLEU 0.194 0.156 0.142 0.163 0.085 0.208 0.158
CHRF 0.236 0.156 0.169 0.232 0.063 0.244 0.183
BERTScore 0.263 0.311 0.197 0.325 0.236 0.336 0.278
BARTScore 0.254 0.287 0.201 0.201 0.182 0.317 0.240
COMET 0.476 0.403 0.417 0.432 0.396 0.417 0.424
COMET-KIWI 0.392 0.367 0.354 0.475 0.442 0.395 0.404
LLaMA Single 1B 0.255 0.371 0.286 0.530 0.518 0.320 0.380
3B 0.284 0.366 0.293 0.578 0.526 0.293 0.390
8B 0.363 0.376 0.356 0.632 0.574 0.491 0.465
Ensemble (3B,1B) 0.287 0.371 0.307 0.568 0.538 0.306 0.396
(8B,1B) 0.325 0.382 0.354 0.605 0.575 0.462 0.450
(8B,3B) 0.337 0.379 0.348 0.619 0.574 0.453 0.452
Contrast (3B,1B) 0.338 0.382 0.347 0.599 0.562 0.284 0.419
(8B,1B) 0.392 0.391 0.406 0.641 0.590 0.484 0.484
(8B,3B) 0.383 0.393 0.409 0.639 0.595 0.482 0.483
Qwen Single 0.5B 0.222 0.394 0.294 0.487 0.557 0.347 0.383
3B 0.306 0.413 0.299 0.573 0.594 0.415 0.433
7B 0.326 0.419 0.330 0.600 0.574 0.445 0.449
Ensemble (3B,0.5B) 0.290 0.408 0.315 0.548 0.599 0.421 0.430
(7B,0.5B) 0.301 0.412 0.331 0.567 0.598 0.446 0.443
(7B,3B) 0.333 0.424 0.341 0.599 0.608 0.456 0.460
Contrast (3B,0.5B) 0.342 0.432 0.351 0.590 0.628 0.431 0.463
(7B,0.5B) 0.359 0.435 0.382 0.611 0.628 0.465 0.480
(7B,3B) 0.362 0.439 0.384 0.605 0.629 0.466 0.481
Table 2: Pearson correlation of evaluators with human scores in machine translation. bold represents best overall scores, while underline represents best scores within each model group (Baseline, LLaMA, Qwen). Overall, ContrastScore outperforms single and ensemble methods, as well as baseline metrics for both LLaMA and Qwen families.
@@ 5.2 Bias Analysis
Table 3 presents the likelihood bias scores across different evaluators of the LLaMA family. The results clearly demonstrate that ContrastScore substantially mitigates likelihood bias compared to both single-model and ensemble approaches. Specifically, ContrastScore achieves lower likelihood bias than any of the individual models employed. For instance, using LLaMA (8B, 3B), it reduces likelihood bias by 18.0% and 64.3% compared to the single LLaMA 8B model on QAGS-XSUM and MQM22, respectively.
ContrastScore is also effective in mitigating likelihood bias for the Qwen model family, although the improvements are somewhat smaller than those observed with LLaMA. Notably, ContrastScore consistently outperforms all individual Qwen models in reducing likelihood bias, with the exception of the 0.5B model on MQM23. These findings suggest that ContrastScore is highly effective in diminishing the reliance on the inherent likelihood of individual models. See also Tables 8 and 9 in Appendix for more details and breakdown results.
Settings Machine Translation Summarization
MQM22 MQM23 Q-XSUM SummEval
Single
1B 0.342 0.212 0.382 0.348
3B 0.323 0.245 0.289 0.385
8B 0.297 0.352 0.267 0.381
Ensemble
(3B,1B) 0.215 0.123 0.249 0.308
(8B,1B) 0.180 0.152 0.225 0.326
(8B,3B) 0.222 0.229 0.242 0.359
Contrast
(3B,1B) 0.058 0.026 0.233 0.183
(8B,1B) 0.104 0.134 0.220 0.262
(8B,3B) 0.106 0.137 0.219 0.240
Table 3: Likelihood bias scores for machine translation and summarization tasks across LLaMA model family. The lowest overall bias score is boldfaced.
Settings Machine Translation Summarization
MQM22 MQM23 Q-XSUM SummEval
Single
0.5B 0.341 0.252 0.347 0.376
3B 0.442 0.451 0.349 0.392
7B 0.441 0.463 0.373 0.398
Ensemble
(3B,0.5B) 0.331 0.327 0.282 0.379
(7B,0.5B) 0.286 0.294 0.289 0.393
(7B,3B) 0.358 0.381 0.345 0.369
Contrast
(3B,0.5B) 0.307 0.314 0.294 0.236
(7B,0.5B) 0.287 0.302 0.318 0.296
(7B,3B) 0.272 0.287 0.353 0.329
Table 4: Likelihood bias scores for machine translation and summarization tasks across Qwen model family. The lowest overall bias score is boldfaced.